# MAPS: A Multilingual Benchmark for Agent Performance and Security

### v2

@@ 6 Conclusions
We introduce MAPS, the first multilingual benchmark suite for evaluating agentic AI systems, addressing a critical gap in assessing language-specific performance and safety limitations. By extending four widely used agentic benchmarks—GAIA, SWE-bench, MATH, and ASB—into eleven diverse languages, our suite enables the analysis of agent behavior under multilingual conditions. Constructed through a hybrid translation pipeline and human verification, MAPS ensures high linguistic and structural consistency. Experiments reveal critical gaps in performance and robustness when agents operate in non-English settings, particularly in tasks involving natural language reasoning and safety-critical behavior. These findings underscore the importance of language-aware evaluation and multilingual adaptation, especially for real-world agentic deployments. We view MAPS as a practical foundation for building more inclusive, resilient, and globally reliable agentic AI systems, and invite the community to extend it.
@@ A.1 Reproducibility Resources
We release the full MAPS dataset, which includes translations of four agentic AI benchmarks (GAIA, SWE-bench, Math, and Agent Safety Benchmark) across 12 languages. A verified subset of MAPS, manually reviewed by native speakers, is also included for higher-quality evaluations. Links to both the full dataset and verified subset can be found at: MAPS Benchmark. We also provide the code for our hybrid translation pipeline, including prompt design and format-preserving techniques, available at: Hybrid Translation Implementation. All resources are released under the MIT License, allowing unrestricted use, modification, and distribution for academic and commercial purposes, provided proper attribution is given.
A.2 Agentic AI Benchmark Comparison
@@
Agent-SafetyBench (Zhang et al., 2024) Security Semi-agentic Jailbreaks, data leakage, unsafe tool use 2,200 Close
@@ A.3 Hybrid Translation Pipeline Implementation Additional Information
For the translation pipeline, we used Python 3.11 alongside a hybrid setup comprising Google’s Neural Machine Translation (NMT) API, Cohere Command-A, and GPT-4o. To support high-quality multilingual alignment, we crafted dedicated system prompts for each stage of the pipeline—translation verification, semantic enhancement, and structural integrity checks—tailored to the specific requirements of each domain. For example, GAIA tasks involve real-world decision-making, where preserving subtle intent and contextual detail across languages is critical. In contrast, MATH tasks require precise handling of symbolic notation and mathematical terminology, necessitating strict preservation of LaTeX expressions and domain-appropriate language. An example of a task-specific system prompt is provided in Fig. 6. The full translation pipeline implementation, including prompt templates and evaluation logic, is available  [^4] https://tinyurl.com/hybrid-translation-pipeline.
To evaluate translation quality across different configurations, we employed LLM-based evaluation using a structured prompt. This approach follows the framework introduced by Jiang et al.(Kocmi and Federmann, 2023), who show that LLMs can reliably assess translation quality when guided by clear criteria. We adapted and extended their methodology to include four dimensions tailored to our benchmark context (those metrics are similar to the ones we define in the manual verification experiment): (1) adequacy—semantic fidelity to the original task, (2) fluency—naturalness and grammaticality in the target language, (3) formatting accuracy—preservation of structural elements such as LaTeX, code, or tool references, and (4) answerability—whether the translation remains usable and complete. The evaluation prompt used is provided in Fig.7. Unlike traditional evaluation metrics such as COMET (Rei et al., 2020), which rely on the same-language reference translations and embedding-based comparisons, our agentic tasks often lack ground truth references and span cross-language instruction. In such settings, reference-based metrics become inapplicable, whereas LLM-based zero-shot evaluation enables context-sensitive, reference-free scoring.
@@ A.4.2 Translation Verification Extended Results
To ensure the validity of our multilingual benchmark, human verifiers manually reviewed approximately 25% of the translated data—corresponding to over 2,000 samples—across all tasks and languages. In the main manuscript, we reported the overall scores for adequacy, fluency, formatting accuracy, and the error rate derived from answerability. In Tables 6, 7, 8 and 9 we present the full breakdown of results for each individual dataset and in Table 10 we conclude with the manual verification average results across all datasets. Across all four benchmark datasets, native verifiers’ ratings reveal high translation quality, with average adequacy (4.47), fluency (4.60), and formatting accuracy (4.76) on a 1–5 scale. Notably, answerability remained consistently high (mean: 0.94), suggesting that most translations retained functional fidelity. Formatting consistency was strongest for ASB and GAIA, while fluency slightly declined for Arabic and Japanese, correlating with minor adequacy drops. Additionally, Fig. 8 presents one successful and one failed translation example, as annotated by Japanese and Korean evaluators, respectively. These results confirm that closed-form multilingual alignment is achievable across diverse tasks without compromising validity.
@@ A.4.3 Statistical sufficiency of the 25% manual audit
Thus, even in the worst case, the corpus-level answerability remains above 92 %.
@@ A.4.4 Language Variants Used
Figure 10: Relative performance differences from English for each language, broken down by dataset. Negative values indicate a drop in performance compared to English, while positive values (notably in ASB) represent increased vulnerability. The trend highlights how multilingual effects vary by language and task type.
A.5 Experimental Settings Additional Information
@@
A.5.1 Agents Configuration Additional Information
@@
A.6 Results Additional Information
@@
This section supplements the main manuscript with detailed quantitative and qualitative results. While the main paper summarizes overall trends and multilingual effects, Tables 1114 provide per-language scores, averages, and standard deviations for each benchmark. Fig. 9 provides a multilingual qualitative failure example of Hindi language in GAIA, including a comparison of main steps in English and Hindi. Fig.  10 provides a language-wise breakdown of relative performance degradation across datasets. From the Figure, we can see that there is no clear correlation between multilingual security robustness (ASB) and multilingual performance degradation. This disconnect is especially clear in real-world, language-heavy tasks like GAIA, where performance drops sharply, while structured tasks like SWE-bench and MATH remain largely unaffected. This highlights that multilingual security alignment does not directly track with multilingual task accuracy, notably in language-rich agentic tasks.

### v3

@@ 6 Conclusions
We introduce MAPS, the first multilingual benchmark suite for agentic AI systems, addressing a critical gap in assessing language-specific performance and security limitations. By extending four widely used agentic benchmarks—GAIA, SWE-Bench, MATH, and ASB—into eleven diverse languages, our suite enables the analysis of agent behavior under multilingual conditions. Constructed through a hybrid translation pipeline and human verification, MAPS ensures high linguistic and structural consistency. Experiments reveal critical gaps when agents operate in non-English settings, particularly in tasks involving natural language reasoning and safety-critical behavior. These findings underscore the importance of language-aware evaluation and multilingual adaptation, especially for real-world agentic deployments. We view MAPS as a practical foundation for building more inclusive, resilient, and globally reliable agentic AI systems, and invite the community to extend it.
7 Limitations
MAPS introduces the first multilingual suite for evaluating agentic AI systems, spanning four domains and eleven languages. While this represents a strong foundation, expanding into additional domains—such as healthcare, or legal reasoning, and into extremely low-resource languages (e.g., Amharic, Uyghur) would broaden its applicability and strengthen its diagnostic power.
A more specific limitation arises from our decision to vary only the input language, while keeping the agent environments fixed in English. This design reflects the dominant real-world regime where users interact in multiple languages, but the underlying resources—documentation, codebases, schemas, or web content—remain primarily English. While this setup simplifies comparison and isolates linguistic understanding as the primary variable, it does not capture the full complexity of localized ecosystems. Fully non-English environments could introduce additional challenges stemming from tool compatibility, dataset structure, or culturally specific context. Nevertheless, our fixed-environment design offers clear advantages: it ensures cross-language comparability, maintains consistent evidence and outputs, and provides a clean diagnostic for language-conditioned reasoning. We therefore view MAPS as an essential first step toward localized multilingual evaluation—an upper bound on agent robustness that establishes a stable foundation for future extensions.
@@ A.1 Reproducibility Resources
We release the full MAPS dataset, which includes translations of four agentic AI benchmarks (GAIA, SWE-Bench, Math, and Agent Security Benchmark) across 12 languages. A verified subset of MAPS, manually reviewed by native speakers, is also included for higher-quality evaluations. Links to both the full dataset and verified subset can be found at: https://huggingface.co/datasets/Fujitsu-FRE/MAPS. We also provide the code for our hybrid translation pipeline, including prompt design and format-preserving techniques, available at: https://github.com/omerhof-fujitsu/hybrid_translation_demo/.
@@
A.2 Agentic AI Benchmark Comparison
The columns in Tables 5 and 6 summarize key properties of each benchmark as follows. Evaluation Objective specifies what the benchmark is primarily designed to measure. Performance benchmarks evaluate the agent’s accuracy to fulfill the given tasks, whereas Security benchmarks probe robustness to adversarial inputs. Task Scope describes the degree of freedom given to the agent. Fully agentic benchmarks specify only the task and success criterion, allowing arbitrary agent architectures and strategies, whereas semi-agentic benchmarks provide a partially fixed agentic scaffold (e.g., code templates or tool interfaces) and mainly vary the underlying LLM or prompting. Domain Coverage indicates the application domains spanned by the benchmark (e.g., software engineering, mathematics, web navigation, security). of tasks reports the number of evaluation instances included in the benchmark. Answer Format specifies how solutions are expressed and evaluated. Closed formats require producing a result in a predefined, structured form with a definitive correct answer, whereas open formats allow free-form outputs for which multiple solutions may be valid. Ground Truth Available indicates whether each task has an explicit reference solution. Multilingual denotes whether the benchmark natively supports evaluation in multiple languages, as opposed to being designed only for English inputs.
@@ A.2 Agentic AI Benchmark Comparison
Below, we provide additional details regarding each benchmark:
@@ A.3 Hybrid Translation Pipeline Implementation Additional Information
For the translation pipeline, we used Python 3.11 alongside a hybrid setup comprising Google’s Neural Machine Translation (NMT) API, Cohere Command-A, and GPT-4o. To support high-quality multilingual alignment, we crafted dedicated system prompts for each stage of the pipeline—translation verification, semantic enhancement, and structural integrity checks—tailored to the specific requirements of each domain. For example, GAIA tasks involve real-world decision-making, where preserving subtle intent and contextual detail across languages is critical. In contrast, MATH tasks require precise handling of symbolic notation and mathematical terminology, necessitating strict preservation of LaTeX expressions and domain-appropriate language. An example of a task-specific system prompt is provided in Fig. 6.
To evaluate translation quality across different configurations, we employed LLM-based evaluation using a structured prompt. This approach follows the framework introduced by Kocmi and Federmann (2023), who show that LLMs can reliably assess translation quality when guided by clear criteria. We adapted and extended their methodology to include four dimensions tailored to our benchmark context (those metrics are similar to the ones we define in the manual verification experiment): (1) adequacy—semantic fidelity to the original task, (2) fluency—naturalness and grammaticality in the target language, (3) formatting accuracy—preservation of structural elements such as LaTeX, code, or tool references, and (4) answerability—whether the translation remains usable and complete. The evaluation prompt used is provided in Fig.7. Unlike traditional evaluation metrics such as COMET (Rei et al., 2020), which rely on the same-language reference translations and embedding-based comparisons, our agentic tasks often lack ground truth references and span cross-language instruction. In such settings, reference-based metrics become inapplicable, whereas LLM-based zero-shot evaluation enables context-sensitive, reference-free scoring.
@@ A.4.2 Translation Verification Extended Results
To ensure the validity of our multilingual benchmark, human verifiers manually reviewed approximately 25% of the translated data—corresponding to over 2,000 samples—across all tasks and languages. In the main manuscript, we reported the overall scores for adequacy, fluency, formatting accuracy, and the error rate derived from answerability. In Tables 10, 11, 12 and 13 we present the full breakdown of results for each individual dataset and in Table 14 we conclude with the manual verification average results across all datasets. Across all four benchmark datasets, native verifiers’ ratings reveal high translation quality, with average adequacy (4.47), fluency (4.60), and formatting accuracy (4.76) on a 1–5 scale. Notably, answerability remained consistently high (mean: 0.94), suggesting that most translations retained functional fidelity. Formatting consistency was strongest for ASB and GAIA, while fluency slightly declined for Arabic and Japanese, correlating with minor adequacy drops. Additionally, Fig. 8 presents one successful and one failed translation example, as annotated by Japanese and Korean evaluators, respectively. These results confirm that closed-form multilingual alignment is achievable across diverse tasks without compromising validity.
@@ A.4.3 Statistical sufficiency of the 25% manual audit
Thus, even in the worst case, the corpus‑level answerability remains above 92 %.
@@
Feature Mean AUC Interpretation
Answerability 0.62 Mild positive relation to accuracy
Formatting 0.55 Very weak positive relation
Adequacy 0.51 $\approx 0.5$, non-predictive
Fluency 0.49 $\approx 0.5$, non-predictive
Table 9: Correlation between human-rated translation features and benchmark accuracy across non-English languages. Answerability shows the strongest positive relationship with correctness, while adequacy, fluency, and formatting hover near chance level, indicating limited predictive power.
@@ A.4.4 Language Variants Used
A.5 Translation-Quality Correlation Analysis.
To quantify how human-rated translation features relate to task accuracy, we conducted a within-task correlation analysis across all non-English languages. For each task, we computed the area under the ROC curve (AUC) between the binary correctness label (“correct answer”) and each translation-quality feature: adequacy, fluency, formatting, and answerability. Averaging these per-task AUCs controls for task difficulty and provides a robust measure of how predictive each feature is of successful task transfer.
These results confirm that while adequacy, fluency, and formatting ensure overall translation fidelity, they do not meaningfully predict whether a translated prompt remains solvable. Answerability instead captures this functional dimension—whether sufficient information is preserved for the correct reasoning process to still hold—making it the most informative indicator of translation quality for multilingual benchmark reliability.
@@
Figure 10: Relative performance differences from English for each language, broken down by dataset. Negative values indicate a drop in performance compared to English, while positive values (notably in ASB) represent increased ASR. The trend highlights how Multilingual Effects vary by language and task type.
A.6 Experimental Settings Additional Information
@@
A.6.1 Agents Configuration Additional Information
@@
English Arabic Chinese German Hebrew Hindi Italian Japanese Korean Portuguese (BR) Russian Spanish
MATH-Chat 51.2% 51.8% 53.9% 50.7% 52.9% 44.5% 48.9% 44.7% 42.7% 51.3% 51.8% 54.5%
OpenDeepResearch (ODR) 51.0% 47.2% 48.4% 52.7% 51.5% 47.1% 49.2% 51.0% 48.3% 51.0% 51.2% 48.7%
Table 15: Cross-benchmark evaluation on the MATH benchmark. We compare MATH-Chat, a benchmark-specific agent, with OpenDeepResearch (ODR), our GAIA agent evaluated under the same multilingual setup. Despite architectural and domain differences, ODR achieves nearly identical performance in English (0.1% difference) and exhibits negligible multilingual degradation (0.11%), supporting the conclusion that cross-lingual degradation depends on linguistic load rather than agent design.
English Arabic Chinese German Hebrew Hindi Italian Japanese Korean Portuguese (BR) Russian Spanish
GAIA Benchmark (Accuracy ↑, higher is better)
GPT-O1 47.3% 31.6% 40.8% 35.1% 35.4% 34.0% 33.2% 36.0% 34.5% 37.4% 39.1% 35.0%
GPT-4o 40.1% 29.0% 36.4% 33.9% 29.0% 32.0% 33.3% 31.4% 26.5% 33.9% 32.7% 30.2%
ASB Benchmark (Attack Success Rate ↓, lower is better)
GPT-5 mini 32.7% 42.5% 58.7% 40.0% 40.5% 55.0% 39.2% 56.0% 53.0% 57.0% 39.5% 39.2%
Qwen 2 7B 48.2% 51.6% 44.7% 65.0% 57.1% 63.3% 59.2% 48.9% 57.0% 63.8% 61.5% 57.6%
Table 16: Cross-Model Agentic Stability Evaluation. Comparison of two base LLMs for the same agentic setups across multilingual settings: Qwen 2 7B vs. GPT-5 mini (ASB security benchmark) and GPT-4o vs. GPT-O1 (GAIA performance benchmark). Metrics are reported per language. ASB: Attack Success Rate (↓ lower is better); GAIA: Accuracy (↑ higher is better). Results show that multilingual degradation trends persist independently of the underlying LLM, supporting the findings reported in Section 4.2 in the main paper.
English Arabic Chinese German Hebrew Hindi Italian Japanese Korean Portuguese (BR) Russian Spanish
OpenDeepResearch GPT-O1 (GAIA Benchmark, Accuracy ↑, higher is better)
Original 47.0% 31.0% 40.0% 35.0% 35.0% 31.0% 33.0% 36.0% 34.0% 37.0% 34.0% 35.0%
Agent decision 38.1% 41.8% 39.3% 43.3% 40.0% 37.2% 43.7% 37.6% 41.2% 45.0% 43.1%
Deterministic translation 42.0% 42.8% 43.2% 42.6% 38.8% 40.7% 37.6% 39.5% 46.8% 45.6% 42.6%
Table 17: OpenDeepResearch - ODR (GAIA) Translation Evaluation. Comparing Original, Agent Decision, and Deterministic Translation. Higher is better (accuracy ↑). All values are percentages.
English Arabic Chinese German Hebrew Hindi Italian Japanese Korean Portuguese (BR) Russian Spanish
ASB Agent GPT-5 (ASB Benchmark, Attack Success Rate ↓, lower is better)
Original 32.7% 42.5% 58.7% 40.0% 40.5% 55.0% 39.2% 56.0% 53.0% 57.0% 39.5% 39.2%
Agent decision 45.0% 36.0% 38.7% 39.0% 53.0% 40.0% 56.0% 54.0% 39.0% 38.0% 38.0%
Deterministic translation 42.5% 35.5% 38.7% 48.0% 48.0% 35.5% 49.0% 44.2% 30.5% 37.2% 35.7%
Table 18: ASB Agent (ASB) Translation Evaluation. Comparing Original, Agent Decision, and Deterministic Translation. Lower is better (attack success rate ↓). All values are percentages.
A.7 Results Additional Information
This section supplements the main manuscript with detailed quantitative and qualitative results.
Cross-Benchmark Evaluation (Extended). A key question in our study is whether the observed multilingual degradation reflects a property of specific agent implementations or a more general limitation of agentic reasoning across languages. To test this, we designed a cross-benchmark experiment—evaluating an agent outside its original domain—to see if the multilingual trend persists.
However, such experiments are inherently constrained by the benchmark-overfitting problem: modern agents are typically engineered for narrow domains, with toolchains, prompting strategies, and even reasoning styles optimized for a single benchmark. As noted by Chang et al. (2024); Gioacchini et al. (2024), this coupling makes cross-benchmark evaluation difficult and often misleading, since performance drops may result from tool incompatibility rather than reasoning failure. For example, SWE-Agent’s code-compilation tools are irrelevant to GAIA’s web-retrieval tasks, and GAIA agents lack the fine-grained debugging environment required for SWE.
Despite these challenges, we identified one feasible cross-benchmark scenario. We evaluated OpenDeepResearch (ODR) - our GAIA agent—on the MATH benchmark, where its quantitative reasoning capabilities (via Python/SymPy) remain relevant. ODR was used without modification to its architecture, prompt, or tool configuration. Remarkably, its performance closely matched the benchmark-specific Math-Chat agent (0.1% difference in average accuracy) and exhibited negligible multilingual degradation (0.11%), mirroring Math-Chat’s own behavior. In contrast, ODR experiences a 12% multilingual drop on the language-heavy GAIA benchmark.
These findings suggest that multilingual degradation is not tied to a specific implementation or architecture, but rather to the linguistic load and reasoning demands of the task environment. In language-light benchmarks such as MATH, both specialized and external agents maintain consistent multilingual performance; in contrast, in linguistically complex benchmarks such as GAIA, ODR performance degrades substantially. Full per-language results are provided in Table 15.
Unified-Backbone Evaluation (extended). Table 19 shows the full per-language results for the unified-backbone experiment described in Section 4.2, in which all four agents are re-evaluated using GPT-5-mini as a shared base LLM, while keeping all other components fixed.
Benchmark EN AR ZH DE HE HI IT JA KO PT-BR RU ES
ASB (ASR %, $\downarrow$ is better)
ASB 32.7 42.5 58.7 40.0 40.5 55.0 39.2 56.0 53.0 57.0 39.5 39.2
GAIA (Accuracy %, $\uparrow$ is better)
GAIA 55.3 45.5 46.7 48.8 49.7 43.4 46.3 50.9 43.0 46.1 44.2 43.3
MATH (Accuracy %, $\uparrow$ is better)
MATH 82 77 84 83 83 79 83 81 77 81 81 83
SWE-Bench (Accuracy %, $\uparrow$ is better)
SWE-Bench 32 29 31 33 32 30 34 29 35 33 34 29
Table 19: Full per-language results for the unified-backbone experiment, in which all agents are re-evaluated using GPT-5-mini as a shared base LLM. Despite higher absolute performance, the same pattern persists: language-heavy benchmarks (GAIA, ASB) exhibit substantial multilingual degradation, while language-light benchmarks (MATH, SWE-Bench) remain largely stable.
Translation evaluation results (extended). To assess whether multilingual degradation arises primarily from translation quality or from deeper linguistic factors, we evaluated two translation strategies across representative agent–LLM pairs. For each benchmark, we compared: (1) the original multilingual prompts (direct human or NMT translation), (2) the agent-decision mode, in which the agent autonomously decides whether and how to translate its inputs before reasoning or tool invocation, and (3) a deterministic-translation mode, where all non-English inputs are consistently translated into English prior to inference and then back-translated for evaluation. This design isolates the contribution of translation control from broader task-specific differences.
For ODR (GAIA) (Table 17), translation likewise improves performance across most languages, particularly under deterministic control, yet the gap with English persists. Agent-driven translation helps moderately but remains inconsistent, highlighting that model-internal translation heuristics are insufficient for stable reasoning across languages. Overall, these experiments demonstrate that while translation—whether autonomous or deterministic—can partially mitigate multilingual degradation, it cannot eliminate it. A persistent, language-specific effect remains even after all text is nominally brought into English, pointing to deeper representational and alignment disparities in current LLM architectures.
Results for ASB Agent (ASB) are shown in Table 18. Both translation strategies reduce attack-success rates relative to the original multilingual runs, with the deterministic approach showing the clearest improvement for languages such as Italian, Portuguese (BR), and Chinese. However, neither strategy fully restores English-level robustness: even when prompts are deterministically translated, residual cross-lingual variability remains. This suggests that part of the multilingual security gap stems not only from translation noise but also from underlying linguistic and policy-alignment differences within the model itself.
Main Experiment Result (extended). While the main paper summarizes overall trends and Multilingual Effects, Tables 2023 provide per-language scores, averages, and standard deviations for each benchmark.
Additionally, we have experimented with 2 additional base-models in the GAIA and ASB datasets. Namely, we apply GPT-4o to the OpenDeepResearch agent used in the GAIA dataset and Qwen-2 base model for the ASB agent used in the ASB dataset. Table 16 presents the performance and ASR of both agents with the new base-LLM, respectively. Interestingly, in the evaluation of the Qwen 7B model in ASB, we observe that Chinese and Japanese yield the lowest ASR in the ASB benchmark, indicating the highest robustness to adversarial prompts in these two Asian languages. We attribute this pattern to the choice of backbone: the ASB agent was implemented using the Qwen2 model  (Yang et al., 2024a), which is known for its strong alignment for Chinese and Japanese language tasks. Qwen2 has consistently demonstrated strong performance in Chinese and Japanese-specific LLM benchmarks Rinna Co. (2025), suggesting that alignment to a particular language and strategic backbone selection can significantly enhance resilience against multilingual adversarial prompts.
Multilingual qualitative failure example. Fig. 9 provides a multilingual qualitative failure example of Hindi language in GAIA, including a comparison of the main steps in English and Hindi.
@@ A.7 Results Additional Information
Language-wise breakdown of performance and ASR results. Fig.  10 provides a language-wise breakdown of relative performance degradation across datasets. From the Figure, we can see that there is no clear correlation between multilingual security robustness (ASB) and multilingual performance degradation. This disconnect is especially clear in real-world, language-heavy tasks like GAIA, where performance drops sharply, while structured tasks like SWE-Bench and MATH remain largely unaffected. This highlights that multilingual security alignment does not directly track with multilingual task accuracy, notably in language-rich agentic tasks.